A research study was conducted to determine if children with attention deficit/hyperactivity disorder, ADHD, have smaller brains than children without the condition. She computed a z score of -1.73.

Using a left tailed test of hypothesis, find the p-value, and at a 5% confidence level, determine if the data provides evidence that children with ADHD have smaller brains than the children without the condition.
Answer #1

Given that,

Zscore =-1.73

Given test is left tailed test.

Null and alternative hypothesis,

Ho:\mu=\muo

H1:\mu<\muo

We want to calculate p-value are as follow,

P-value =P(Z<-1.73)

=1- P(Z<1.73)

By using Standard Normal table

P(Z<1.73)=0.9582

Therfore,

P-value =1-0.9582=0.0418

P-value =0.0418

Conclusion:

since, P-value =0.0418 < 0.05

It is concluded that the rejecte the null hypothesis, Therefore, there is sufficient evidence to claim that the children with ADHD have smaller brains than the children without the condition.